OpenStudy (anonymous):

Find the differential dy of the given function. (Use "dx" for dx.) y = xsqrt(9 − x^2) please help

OpenStudy (turingtest):

can you take the derivative of this function?

OpenStudy (anonymous):

yes its \[\frac{ 9-2x ^{2} }{ \sqrt{9-x ^{2}} }\]

OpenStudy (freckles):

\[\frac{dy}{dx}=\frac{9-2x^2}{\sqrt{9-x^2}} \\ dy =\frac{9-2x^2}{\sqrt{9-x^2}} dx\]

OpenStudy (freckles):

oh this was 13hrs ago lol